# How Good Are We? Evaluating Cell AI Foundation Models in Kidney Pathology with Human-in-the-Loop Enrichment

### v1

@@ Abstract
Training AI foundation models has emerged as a promising large-scale learning approach for addressing real-world healthcare challenges, including digital pathology. While many of these models have been developed for tasks like disease diagnosis and tissue quantification using extensive and diverse training datasets, their readiness for deployment on some arguably simplest tasks, such as nuclei segmentation within a single organ (e.g., the kidney), remains uncertain. This paper seeks to answer this key question, “How good are we?”, by thoroughly evaluating the performance of recent cell foundation models on a curated multi-center, multi-disease, and multi-species external testing dataset. Additionally, we tackle a more challenging question, “How can we improve?”, by developing and assessing human-in-the-loop data enrichment strategies aimed at enhancing model performance while minimizing the reliance on pixel-level human annotation. To address the first question, we curated a multicenter, multidisease, and multispecies dataset consisting of 2,542 kidney whole slide images (WSIs). Three state-of-the-art (SOTA) cell foundation models—Cellpose, StarDist, and CellViT—were selected for evaluation. To tackle the second question, we explored data enrichment algorithms by distilling predictions from the different foundation models with a human-in-the-loop framework, aiming to further enhance foundation model performance with minimal human efforts. Our experimental results showed that all three foundation models improved over their baselines with model fine-tuning with enriched data. Interestingly, the baseline model with the highest F1 score does not yield the best segmentation outcomes after fine-tuning. This study establishes a benchmark for the development and deployment of cell vision foundation models tailored for real-world data applications.
keywords:
Foundation Models, Segmentation Performance, Active Learning, Ensemble Learning, Domain Adaptation, Kidney
@@
1 Introduction
AI foundation models trained on massive, diverse datasets are widely applied across numerous fields, including healthcare [1, 2]. Their versatility enables these models to tackle a variety of downstream tasks, with one of the most prominent applications being digital pathology [3, 4, 5, 6]. Among many tasks in digital pathology, cell instance segmentation is often serving as the initial step in extracting biological signals crucial for accurate disease diagnosis and treatment planning [7, 8, 9, 10, 11, 12]. The accuracy of the cell or cell nuclei segmentation forms the foundation for various subsequent biological or medical analyses, including cell type classification [13], specific cell counting [14], and cell phenotype analysis [15]. It is also considered as the stepping stone for whole slide image (WSI) analysis in any biological and biomedical applications [16].
In recent years, several cell foundation models [17, 18, 19, 20] trained on large and diverse datasets—encompassing various cell types, imaging techniques (e.g., fluorescence, brightfield, phase contrast), and experimental conditions—have demonstrated promising results. While many of these models have been developed for tasks such as disease diagnosis and tissue quantification, their readiness for deployment on some of the arguably simpler tasks, such as nuclei segmentation within a single organ (e.g., the kidney), remains uncertain. This leads us to pose the first question: Is nuclei segmentation on whole slide images (WSIs) within a single organ, like the kidney, a solved problem using current cell foundation models in histopathology?
To answer the preceding question, we have chosen the kidney as the evaluation organ for this study due to its diverse cell types. According to [21], the kidney comprises at least 16 specialized epithelial cell types, along with various endothelial, immune, and interstitial cells. Additionally, kidney whole slide images and their primary staining type, Periodic Acid-Schiff (PAS), are underrepresented in the training of current cell nuclei foundation models, underscoring a gap in the research field. We evaluated the performance of current cell nuclei foundation models in kidney pathology within a multi-center, multi-staining, multi-species setting. As shown in Fig. 1a, b, and c, we construct a diverse evaluation dataset that includes kidney nuclei data from 2,542 kidney whole slide images (WSIs) sourced from humans and rodents across both public and in-house datasets. To our knowledge, the scale of this study’s kidney WSIs surpasses all publicly available labeled nuclei datasets that include the kidney, as illustrated in Fig. 1a. We conducted a comparative analysis of three widely used SOTA cell nuclei foundation models—Cellpose, StarDist, and CellViT—for kidney cell nuclei instance segmentation. Model performance was evaluated by collectively rating each model’s prediction mask as “good,” “medium,” or “bad.” Specifically, “good” predictions were defined as those capturing approximately 90% of the nuclei in an image patch, while “medium” predictions captured between 50% and 90% of the nuclei. Predictions rated as “bad” captured less than 50% of the nuclei. This quantitative definition clarifies the criteria for each rating type, based on a standard established by a renal pathologist at Vanderbilt University Medical Center with over 20 years of experience. Our evaluation results indicated that a performance gap still exists in general nuclei segmentation for kidney pathology. Among the evaluated models, CellViT exhibited the best performance in segmenting kidney nuclei, achieving 63% of “good” ratings on the evaluation dataset.
Then, a natural follow-up question is: can we further enhance the performance of these trained cell foundation models? To tackle this, we performed data enrichment by ensembling predictions from three foundation models within a human-in-the-loop (HITL) design. Specifically, the “bad” image patches, manually corrected by pathologists, incorporated the domain knowledge required by the models, while the “good” predictions distilled insights from various foundation models. We assessed the model’s performance gains from three data enrichment strategies: using “good” image patches only, using “bad” image patches only, and using a combination of both. Our experimental results demonstrated that foundation models can mutually enhance their segmentation capabilities through data enrichment strategies, with StarDist achieving the highest F1 score of 0.82. The baseline model with the highest F1 score (CellViT) does not yield the best segmentation outcomes after fine-tuning. The combination of “good” and “bad” image patches proved to be the most effective strategy. Ultimately, these improved models can facilitate more efficient workflows in annotation software, such as QuPath [22], for kidney pathology.
2 Related work
2.1 Image Processing-Based Nuclei Instance Segmentation
A key challenge in nuclei instance segmentation is separating overlapping nuclei [23, 24, 25, 26, 27, 28, 29, 30]. Solutions to this problem range from traditional image processing techniques to deep learning approaches. Traditional methods rely on image processing for feature extraction, using intensity, shape, texture, and morphological patterns to identify nuclei. For instance, traditional methods such as marker-controlled watershed segmentation have been widely adopted for their effectiveness in separating overlapping or touching nuclei by utilizing predefined markers and adaptive thresholding techniques [23, 26]. However, these methods often struggle with over segmentation or under segmentation when dealing with highly heterogeneous or densely clustered nuclei. To address these limitations, region-based active contour models and shape-prior integration have been developed, which incorporate geometric shape information to improve boundary delineation and handle object occlusions [28]. Additionally, multi-pass segmentation strategies, like the Multi-Pass Fast Watershed (MPFW) method, have demonstrated superior performance by employing iterative segmentation passes and combining gradient-based and shape-based techniques, making them particularly effective for complex cytological samples such as cervical cell images [25]. However, these image processing-based methods remain heavily reliant on hand-crafted features, requiring domain expertise and being sensitive to hyperparameters [31, 32].
2.2 Deep Learning-Based Nuclei Instance Segmentation
To address the limitations of hand-crafted feature-based methods, deep learning (DL) has enabled automatic feature extraction. As outlined in [33], current DL approaches for nuclei instance segmentation can be broadly categorized into two-stage and one-stage approaches. Two-stage methods, such as Mask-RCNN [34], first detect and localize cell nuclei using bounding boxes before refining the segmentation in a subsequent stage. Similarly, [35] uses the first stage to generate nuclei proposals and the second stage to refine nuclei boundaries. However, two-stage methods often suffer from non-standardized training and high computational costs. Additionally, overlapping nuclei can lead to incorrect segmentation being passed from the first to the second stage, requiring further postprocessing.
In contrast, one-stage methods include DL network and postprocessing into an end to end training and inference framework. For example, Micro-Net [36] modifies the U-Net architecture [37] by incorporating multi-resolution input images, enabling better performance for nuclei of different sizes. The Dist model [38] introduces an additional decoder branch to predict distance maps from nucleus boundaries to their center of mass, aiding watershed-based postprocessing. HoVer-Net [31], a SOTA method, leverages horizontal and vertical distance maps to detect nuclei edges through the Sobel operator. StarDist [18] and CPP-Net [32] use a polygon-based approach, which both show comparable results with HoVer-Net. Cellpose [17] incorporates a U-Net architecture combined with gradient flow tracking to create an end-to-end model for nuclei instance segmentation. Similarly, network architecture in CellViT [19] uses a U-shaped vision transformer (ViT) [39] combined with HoVer-Net postprocessing. Overall, one-stage models offer a balance between accuracy and efficiency, making them more suitable for practical deployment where computational resources are limited. Leveraging the ease of standardization and training offered by one-stage methods, this work also incorporates foundation models from this branch.
2.3 Human-in-the-loop
Another key factor in improving nuclei instance segmentation is data labeling, as training instance segmentation models require pixel-level annotations for every object in an image [20]. However, creating these annotations can be expensive ($10^{-2}\text{ to }10^{0}$ USD/label) [9, 40, 41], which highlights the need to reduce the marginal cost of labeling. A recent improvement in labeling involves a multi-phase human-in-the-loop (HITL) process, where pathologists refine model errors to train foundation models sequentially [42, 9, 41]. Once the model achieves human-level performance, it is used to fully automate annotation of new images without further human correction. For example, [41] uses a two-phase HITL strategy to sequentially train models for cell segmentation and tracking by alternating between model predictions and human corrections. [9] uses a three-phase HITL strategy to build a large training dataset through crowdsourced correction and expert validation. Cellpose 2.0  [42] implements a flexible HITL pipeline for fine-tuning pretrained models with minimal new annotations. This process reduces the annotation load to 100–200 ROIs (region of interest) while still achieving high-quality segmentations. However, these approaches do not leverage multiple foundation models with HITL. Motivated by this limitation, in our work, we utilized multiple foundation models with a HITL design to scale up the training dataset and further enhance the performance of the foundation models.
@@ 2.3 Human-in-the-loop
Figure 2: Human-in-the-loop (HITL) Data Enrichment Design. The upper panel shows the inference and curation of prediction masks from three foundation models (Cellpose, StarDist, and CellViT) in this study. First, kidney nuclei instance segmentation was performed on the evaluation dataset using three cell foundation models. Model performance was evaluated by rating each prediction mask as “good,” “medium,” or “bad” according to criteria from a renal pathologist. “Good” predictions captured approximately 90% of the nuclei in a patch, “bad” predictions captured less than 50%, and the rest were classified as “medium.” We used this rating system to both qualitatively and quantitatively evaluate and categorize each model’s predictions within our dataset. Each rating reflects the segmentation quality of the predicted mask for an image patch. Based on the results of the image patch curation, the lower panel illustrates the data enrichment strategy that utilizes these curation outcomes to enhance model performance through continuous fine-tuning. Specifically, to enrich the training dataset while minimizing pixel-level annotation, we ensembled “good” prediction masks from multiple foundation models (curation cost: 10 seconds per image) and actively corrected strongly agreed-upon “bad” samples (annotation cost: 20 minutes per image).
3 Method
This section details the performance evaluation and HITL data enrichment framework (as shown in Fig. 2) used in this study. First, we describe the construction of a diverse, large-scale dataset for evaluating foundation models. Then, we present the HITL design that utilizes evaluations of multiple foundation models to enrich the training dataset. Lastly, the continuous fine-tuning of foundation models using the enriched dataset.
3.1 Curate a Diverse Large-Scale Dataset
To provide a comprehensive assessment of their performance in segmenting kidney nuclei, we constructed a diverse evaluation dataset. Our dataset comprises both public and private kidney data, with a total of 2,542 whole-slide images (WSIs). As illustrated in Fig.1b, 57% (1,449 WSIs) come from publicly available sources, including the Kidney Precision Medicine Project (KPMP)[43], NEPTUNE [44], and HUBMAP [45, 46], while the remaining 43% (1,093 WSIs) originate from an internal collection at Vanderbilt University Medical Center. To increase the diversity of our dataset, we incorporated WSIs from both human and rodent samples. As depicted in Fig. 1c, these WSIs were stained using Hematoxylin and Eosin (H&E), Periodic acid-Schiff methenamine (PASM), and Periodic acid-Schiff (PAS), with PAS being widely used in kidney pathology but less frequently in other organs. We randomly extracted four 512$\times$512-pixel image patches from each kidney WSI at 40$\times$ magnification. Patches that were contaminated, of low imaging quality, from dead tissue, or with incorrect staining methods were discarded. This process resulted in an evaluation dataset of 8,818 image patches.
@@ 3.2 Data Enrichment with Multiple Foundation Models and HITL
As shown in Fig. 2a, nuclei instance segmentation was performed on these kidney nuclei image patches using three cell foundation models: Cellpose [17], StarDist [18], and CellViT [19]. Then, instead of directly correcting their predicted instance masks, we use a rating-based system to evaluate the predictions from the three foundation models (as shown in Fig. 2b). This curation process involves two rounds of input from human experts. In the first round of image patch curation, two pathologist-trained students evaluated the predicted masks, categorizing them as “good,” “medium,” or “bad” based on a rating standard set by a renal pathologist with 20 years of experience. “Good” predictions captured approximately 90% of the nuclei in a patch, “bad” predictions captured less than 50%, and the rest were classified as “medium.” In the second round, the experienced renal pathologist manually reviewed and validated the uncertain samples from the first round. We used this rating system to both qualitatively and quantitatively evaluate each model’s predictions within our dataset. Each rating reflects the segmentation quality of the predicted mask for an image patch. In this study, employing multiple foundation models reduces bias in image patch curation by having each patch evaluated by three models trained with different architectures and settings. This approach offers diverse perspectives, helping to identify image patches that highlight gaps between general and specific (e.g., kidney) domain performance. As a result, curated predictions from these models can be used in the next stage of the HITL process to further refine their performance in segmenting nuclei in kidney pathology.
In this work, we leveraged the image patch curation results to reduce the pixel-level labeling effort in the HITL design. Specifically, we directly utilized the “good” samples (shown as Fig. 2d) for the next round of model refinement. Including these samples and their predictions as image and pseudo-label pairs enhances the training dataset at scale. Additionally, for each foundation model, the “good” samples reflect its strength and pretrained knowledge e(curation cost: 10 seconds per image patch). Pathologists manually corrected or annotated a small set of these “bad” image patches, which can be used for the next round of training for the foundation models (annotation cost: 10 seconds per image patch).
@@ 3.3 Data-Enriched Foundation Models Fine-Tuning
Benefiting from the data enrichment provided by multiple foundation models’ curation, the models can be fine-tuned using either “good” image patches (foundation model-generated pseudo labels), “bad” image patches (pathologist-corrected), or both, to enhance kidney nuclei segmentation performance. In this section, we first introduce the three foundation models used for image patch curation and continuous fine-tuning. A key consideration in fine-tuning, particularly when using both “good” and “bad” image patches, is the imbalance between these two classes. To address this, we applied a customized weighted oversampling method.
3.3.1 Cell Foundation Models
@@
  1. 1.
    Cellpose: Cellpose [17] is a generalist segmentation model that utilizes a U-Net backbone to predict the horizontal and vertical gradients of topological maps, as well as a binary map. It was trained on a diverse dataset containing over 70,000 segmented objects from various microscopy modalities, including fluorescence, brightfield, and other specialized types. The training dataset primarily consists of fluorescently labeled cytoplasm, with or without an additional nuclear channel. Additionally, 30 H&E images from MoNuSeg [47] were included in the nucleus dataset. These H&E-stained histology images contain many small nuclei, and their polarity was inverted to make them resemble fluorescence images more closely. This variety of image sources, covering cells with diverse morphologies and modalities, allowed Cellpose to generalize well across a broad range of cell types without requiring retraining or parameter adjustments. To segment individual cells, Cellpose employs a gradient tracking algorithm that clusters pixels based on their convergence to a common center, enabling precise delineation of cellular boundaries even in challenging cases with dense packing or complex morphologies.
  2. 2.
@@
  3. 3.
    CellViT: CellViT [19] employs a U-Net shaped hierarchical encoder-decoder Vision Transformer (ViT) backbone, designed specifically for nuclei segmentation and classification in histopathology images. The encoder utilizes pre-trained weights from a ViT trained on 104 million histological images (ViT256)[49], a model that demonstrated superior performance in cancer subtyping and survival prediction tasks. CellViT is further trained on the PanNuke dataset [50], which includes 189,744 annotated nuclei across 19 tissue types, grouped into five clinically relevant classes: neoplastic, inflammatory, epithelial, dead, and connective. The images in this dataset are captured at 40$\times$ magnification with a resolution of 0.25 µm/pixel, making it a challenging benchmark due to its diversity and class imbalance. CellViT’s post-processing pipeline follows the HoVer-Net methodology [31], using gradient maps of horizontal and vertical distances for accurate boundary delineation. Additionally, the network benefits from extensive data augmentation techniques and transfer learning strategies to overcome the scarcity of annotated medical data, achieving SOTA performance on both the PanNuke and MoNuSeg datasets.

### v2

# Evaluating Cell AI Foundation Models in Kidney Pathology with Human-in-the-Loop Enrichment

@@ Abstract
Background Large-scale artificial intelligence foundation models have emerged as promising tools for addressing healthcare challenges, including digital pathology. While many have been developed for complex tasks such as disease diagnosis and tissue quantification using extensive and diverse datasets, their readiness for seemingly simpler tasks, such as nuclei segmentation within a single organ (for example, the kidney), remains unclear. This study answers two questions: How good are current cell foundation models? and How can we improve them?
Methods We curated a multi-center, multi-disease, and multi-species dataset sampled from 2,542 kidney whole slide images. Three state-of-the-art cell foundation models—Cellpose, StarDist, and CellViT—were evaluated. To enhance performance, we developed a human-in-the-loop strategy that distilled multi-model predictions, improving data quality while reducing reliance on pixel-level annotation. Fine-tuning was performed using the enriched datasets, and segmentation performance was quantitatively assessed.
Results Here we show that cell nuclei segmentation in kidney pathology still requires improvement with more organ-targeted foundation models. Among the evaluated models, CellViT achieves the highest baseline performance, with an F1 score of 0.78. Fine-tuning with enriched data improves all three models, with StarDist achieving the highest F1 score of 0.82. The combination of the foundation model–generated pseudo-labels and a subset of pathologist-corrected “hard” patches yields consistent performance gains across all models.
Conclusions This study establishes a benchmark for the development and deployment of cell AI foundation models tailored to real-world data. The proposed framework, which leverages foundation models with reduced expert annotation, supports more efficient workflows in clinical pathology.
Plain language summary
The rise of digital pathology has transformed traditional histology slides into vast collections of high-resolution images, enabling medical research on a much larger scale. However, analysing this data remains challenging. Foundation models—advanced AI systems trained on diverse datasets—offer a promising solution, but their ability to perform simpler yet essential tasks, such as identifying cell nuclei in kidney tissue, is unclear. We evaluated three leading models on a large, curated kidney image dataset and found that cell nuclei segmentation in kidney pathology still requires improvement with more organ-targeted foundation models. To enhance performance, we introduced a “human-in-the-loop” approach that combines multiple foundation models with expert labeling of only the most difficult cases, improving accuracy, reducing manual labeling, and enabling more efficient pathology workflows.
1 Introduction
AI foundation models trained on massive, diverse datasets are widely applied across numerous fields, including healthcare [1, 2]. Their versatility enables these models to tackle a variety of downstream tasks, with one of the most prominent applications being digital pathology [3, 4, 5, 6, 7]. In this field, cell instance segmentation serves as a critical first step for extracting biologically meaningful information used in diagnosis, treatment planning, and research [8, 9, 10, 11, 12, 13]. Accurate cell or nuclei segmentation underpins downstream tasks such as cell type classification [14], cell counting [15], spatial transcriptomics [16], and phenotype analysis [17], and is essential for whole slide image (WSI) analysis [18].
The task of segmenting nuclei has evolved considerably over the past two decades. Early solutions relied on image processing techniques such as marker-controlled watershed segmentation [19, 20], active contour models [21], and multi-pass watershed methods [22]. While effective in specific contexts, these methods depended on hand-crafted features and were sensitive to parameter tuning, making them less robust for heterogeneous or overlapping nuclei. The advent of deep learning (DL) addressed many of these limitations by replacing manual feature design with learned representations. Two-stage DL approaches, such as Mask R-CNN [23], first localize nuclei before refining their boundaries, whereas one-stage methods integrate detection and segmentation into a unified process. Notable one-stage models include Micro-Net [24], HoVer-Net [25], StarDist [26], CPP-Net [27], Cellpose [28], and CellViT [29], which vary in architectural design but share the advantages of efficiency and standardization.
More recently, foundation models for cell segmentation [28, 26, 29, 30] have emerged, which are trained on large and diverse datasets encompassing multiple cell types, microscopy techniques, and experimental conditions. These models have demonstrated impressive cross-domain generalization, enabling strong performance in unfamiliar settings. However, despite their success in complex tasks such as disease diagnosis and tissue quantification, it remains unclear whether they can generalize to simpler yet essential tasks, such as cell nuclei segmentation within a single organ. This is a relevant and underexplored question in kidney pathology, where WSIs feature high cellular diversity—at least 16 specialized epithelial types along with endothelial, immune, and interstitial cells [31]—and where common stains such as Periodic Acid–Schiff (PAS) are rarely represented in training datasets.
Beyond large-scale training and model architecture, the availability of high-quality training data remains a critical bottleneck. Producing pixel-level annotations for instance segmentation is time-consuming and expensive (typically requiring $10^{-2}\text{ to }10^{0}$ hours per image) [10, 32, 33]. Human-in-the-loop (HITL) strategies can mitigate this burden by iteratively refining model predictions with expert feedback [34, 10, 33]. For example, Cellpose 2.0 [34] fine-tunes pretrained models using as few as 100–200 region-of-interest annotations. While effective, most HITL pipelines are built around a single model, leaving the potential benefits of combining multiple foundation models largely unexplored.
In this study, we address these gaps by evaluating three widely used cell AI foundation models—Cellpose [28], StarDist [26], and CellViT [29]—on a diverse evaluation dataset that includes kidney nuclei data sampled from 2,542 kidney WSIs sourced from humans and rodents across both public and in-house datasets. To our knowledge, the scale of this study’s kidney WSIs surpasses all publicly available labeled nuclei datasets that include the kidney. Here we show that cell nuclei segmentation in kidney pathology still requires improvement with more organ-targeted models. To enhance performance, we introduce a HITL data enrichment strategy that combines foundation model predictions with expert corrections for the most challenging cases. Fine-tuning with these enriched datasets improves all three models, with StarDist achieving an F1 score of 0.82. Pairing pseudo-labels with a subset of pathologist-corrected “hard” patches consistently boosts performance, improves segmentation accuracy, and streamlines workflows in research and clinical pathology.
2 Methods
In this section, we first introduce the overall framework of the study, followed by details of its design. Specifically, we describe the construction of a diverse, large-scale dataset for evaluating foundation models, the human-in-the-loop strategy for enriching the training data by leveraging these models, the continual fine-tuning of models with the enriched dataset, and the experimental setup.
2.1 Overall Framework
Fig. 1 illustrates the overall framework of this study. First, we curated a diverse, large-scale evaluation dataset comprising samples from 2,542 kidney whole slide images (Fig. 1a-c). Next, kidney cell nuclei instance segmentation was performed using three state-of-the-art foundation models: Cellpose [28], StarDist [26], and CellViT [29]. Model performance was then assessed, as shown in the bottom panel, through qualitative human evaluation of each prediction mask. Lastly, to address the more challenging task of improving model performance, the Fig. 1 Data Enrichment stage employs a human-in-the-loop framework that integrates foundation model predictions with expert corrections for the most difficult cases into the model’s continual learning process, thereby reducing reliance on pixel-level human annotation.
2.2 Curating a Diverse Large-Scale Dataset
@@
Figure 2: Human-in-the-loop (HITL) Data Enrichment Design. The upper panel shows the inference and curation of prediction masks from three foundation models (Cellpose, StarDist, and CellViT) in this study. (a) First, kidney nuclei instance segmentation was performed on the evaluation dataset using three cell foundation models. (b) Model performance was evaluated by rating each prediction mask as “good,” “medium,” or “bad” according to criteria from a renal pathologist. “Good” predictions captured approximately 90% of the nuclei in a patch, “bad” predictions captured less than 50%, and the rest were classified as “medium.” We used this rating system to both qualitatively and quantitatively evaluate and categorize each model’s predictions within our dataset. (c-e) The lower panel illustrates the data enrichment strategy that utilizes these curation outcomes to enhance model performance through continuous fine-tuning. Specifically, to enrich the training dataset while minimizing pixel-level annotation, we used both pseudo-labeled images from multiple foundation models (termed as “easy”) and “hard” samples that all models failed.
@@ 2.3 Data Enrichment with Multiple Foundation Models and HITL
Rating Image Patches. As shown in Fig. 2a, nuclei instance segmentation was performed on these kidney nuclei image patches using three cell foundation models: Cellpose [28], StarDist [26], and CellViT [29]. Then, instead of directly correcting their predicted instance masks, we use a rating-based system to evaluate the predictions from the three foundation models (as shown in Fig. 2b). The ratings, conducted in a blinded experiment, were performed separately by two pathologist-trained students evaluating model predictions. An experienced renal pathologist subsequently reviewed and validated the samples where the students’ assessments disagreed. As shown in Fig. 3, ratings—categorized as “good,” “medium,” or “bad” based on the criteria set by an expert renal pathologist—defined “good” predictions as capturing approximately 90% of nuclei in a patch, “bad” as capturing less than 50%, and all others as “medium.”
Figure 3: Illustrations of Rating Criteria. This Figure illustrates the rating criteria, showing examples and their ratings for “good”, “medium”, and “bad” categories.
Pseudo-Labels and Hard Data Annotation. In this work, we leveraged the image patch curation results to reduce the pixel-level labeling effort in the HITL design. Specifically, we scaled up the training dataset by directly utilizing the “easy” (pseudo-labeled) samples (shown as Fig. 2d). To bridge the domain gap, we also incorporated a small set of representative “hard” samples (shown as Fig. 2c), manually annotated by pathologists at a cost of roughly 20 minutes per patch.
@@ 2.4 Data-Enriched Foundation Models Fine-Tuning
Leveraging curated data from multiple foundation models, we fine-tune them with “easy” patches (pseudo-labels), “hard” patches (pathologist-corrected), or a combination of both to improve kidney-nuclei segmentation. In this section, we (1) introduce the three foundation models used for image-patch curation and continuous fine-tuning, and (2) address a key fine-tuning challenge: the imbalance between “easy” and “hard” image patches. To mitigate this imbalance, we apply a customized weighted-oversampling method.
2.4.1 Cell AI Foundation Models
@@
Cellpose: Cellpose [28] is a generalist segmentation model that utilizes a U-Net backbone to predict the horizontal and vertical gradients of topological maps, as well as a binary map. Then, a gradient tracking algorithm is employed to enable precise delineation of cellular boundaries. According to [28], Cellpose was trained on highly varied images of cells, containing over 70,000 segmented objects, capturing a wide range of imaging modalities and cellular morphologies. The training dataset includes fluorescent microscopy with cytoplasmic markers and DAPI-stained nuclei, brightfield microscopy, membrane-labeled cells, as well as non-cellular and atypical cellular images from Google search to enhance the model’s generalization beyond conventional biological images. Morphologically, it featured diverse cell types, including for example, neurons with complex dendritic processes, bone marrow-derived macrophages, cultured Drosophila and mouse cortical cells, pancreatic stem cells, cancer cell lines such as U2OS osteosarcoma cells, and histological samples from various human organs (e.g., MoNuSeg [39]). This comprehensive diversity in both imaging types and cell morphologies increases its segmentation ability across a wide range of biological and imaging contexts without requiring retraining or parameter adjustments.
@@
[…]